OpenStudy (anonymous):

A 5.0-kg bucket of water is raised from a well by a rope. If the upward acceleration of the bucket is 3.0 m/s2, find the force exerted by the rope on the bucket.

OpenStudy (anonymous):

mass = 5kg NET acceleration = 3 m/s2 F=ma (an equation to memorize) Force, first you gotta defeat gravity. Then more Force to go up. F = 5kg * (3 m/s2 + 9.8 m/s2) = 64 N